OpenStudy (anonymous):

if you have to write Fourier serie for f(x)=x^2, 0

OpenStudy (anonymous):

or why should i have \[a0=2/\pi \int\limits_{0}^{\pi} x ^{2}dx\]

OpenStudy (anonymous):

So we want the Fourier series for F(x)=x^2, for 0<x<pi. You don't have to calculate it for -pi<x<pi. You can use the domain given to you. So the Fourier series would be \[f(x)=\frac{1}{2}a_o+\sum_{n=1}^{\infty}a_ncos(\frac{n \pi x}{L})+b_nsin(\frac{n \pi x}{L})\] for -L<x<L where \[a_0=\frac{1}{L}\int\limits_{-L}^{L}F(x)dx\] \[a_n=\frac{1}{L}\int\limits_{-L}^{L}F(x)cos(\frac{n \pi x}{L})dx\] \[b_n=\frac{1}{L}\int\limits_{-L}^{L}F(x)sin(\frac{n \pi x}{L})dx\]

OpenStudy (anonymous):

So for F(x)=x^2 0<x<pi, the L value will be pi/2. \[a_0=\frac{1}{\frac{\pi}{2}} \int\limits_{0}^{\pi} x^2 dx\] \[=\frac{2}{\pi} \int\limits_{0}^{\pi} x^2 dx\]

OpenStudy (anonymous):

and if you have -pi<x<pi, you use "normal" formula?